OpenStudy (anonymous):

What are the 3 different ways for solving systems of equations and when would you use each one? Substituting, Graphing, and what else?

OpenStudy (anonymous):

@Brainybeauty

OpenStudy (anonymous):

elimination 6x-4y=-20 -6x-8y=-14 Thats an example (used random numbers not sure if that would work

OpenStudy (anonymous):

that would be solving for y

OpenStudy (anonymous):

The third one is elimination

OpenStudy (anonymous):

Ah.. that was easy! Can you help with some more? How are the algebraic properties used in creating equivalent systems?

OpenStudy (anonymous):

???